# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ 2.3 Dynamical tide induced instability and chaos
For binaries with significant eccentricities, the truly dynamical component of a mode can be excited at a pericenter passage. Such a component oscillates at a mode’s natural frequency (hence phase incoherent with the orbit) and does not decrease in amplitude as the companion moves to the apocenter (assuming dissipation is small). In the next pericenter passage, the existing component of the dynamical tide can interfere with the new tidal kick it receives at the pericenter, hence making this component’s evolution history dependent (as opposed to the equilibrium tide that depends only on the instantaneous orbital configuration). This dynamical component formally modifies the stability threshold according to the toy model in Eq. (6), though its complicated evolution prohibits a simple analytical study as we have done in §2.1 and 2.2. Nonetheless, the hydrodynamical equations are sufficiently generic for arbitrary rotations and orbital configurations. We will use a numerical approach to investigate the associated hydrodynamical stability.
@@
In this Subsection, we will solve directly the phase-space expanded equations of $c_{m,\pm}$, Eqs. (15), so that the Coriolis effect can be accounted for naturally. For the driving terms ($K_{m}$), we will consider both the version expanded to the three-wave order, Eqs. (20) and (21), and the exact nonlinear version derived in Appendix A following the affine model of P. Diener et al. (1995) (see also B. Carter & J. P. Luminet 1983, 1985). The monopolar mode $q_{\rm tr}$ is treated fully dynamically, where a differential equation of $\ddot{q}_{\rm tr}$ is solved instead of reducing $q_{\rm tr}$ under the static limit with Eqs. (A31). Note that the affine model in P. Diener et al. (1995) allows for generic values of polytropic index $n$ and adiabatic index $\Gamma$. Thus, our numerical experiments will consider both a homogeneous ellipsoid with $(n,\ \Gamma)=(0,\ 20)$ as before222 Numerically, $\Gamma\to\infty$ is hard to implement. We therefore use a finite $\Gamma=20\gg 1$ as an approximation. , and a polytrope with $(n,\ \Gamma)=(1,\ 2)$ that serves as a good approximation of a buoyancy-neutral Jovian planet. However, a caveat for the $n=1$ model is that the starting point of the affine model (Eq. 12) is not self-consistently satisfied as the Lagrangian displacement $\boldsymbol{\xi}$ is no longer linear in position. We fix the companion’s mass to be $M^{\prime}=10^{3}$ and treat it as a point particle. When computing the orbital motion, we ignore tidal back-reaction and simply model it as Keplerian, with
@@
The initial conditions are set at the apocenter, with the modes computed with their linear equilibrium tide solution. The set of differential equations is evolved using the solve_ivp function in SciPy (P. Virtanen et al., 2020) with an explicit Runge-Kutta method of order 8 (E. Hairer et al., 1993).
@@
In the first experiment shown in Fig. 3, we validate the results of §2.2 by considering the numerical evolution of $|q_{2}|=|c_{2,+}+c_{-2,+}^{\ast}|$ for a non-rotating ($p=0$), homogeneous ($n=0$) star in a moderately eccentric orbit with $e=0.3$. We evolve the system using both the exact nonlinear equations (left panel) and the three-wave-expanded ones (right panel). Qualitatively, the two sets of equations yield similar results, with the exact version allowing the runaway to occur at a slightly further separation $D_{p}/R_{t}\simeq 2.72$. This is consistent with Fig. 1, which shows that the three-wave equations slightly underestimates $D_{p}^{(\rm th)}\propto(1/\epsilon_{\rm max})^{1/3}$.
@@
The accumulation of the dynamical tide becomes more significant in systems with higher eccentricities, with one such example shown in Fig. 4 for a donor moving in a highly eccentric ($e=0.9$) orbit. We still consider a homogeneous donor but vary its rotation rates (left for the non-rotating case and right for the pseudo-synchronized case). Both cases are evolved using the three-wave equations of motion. As the hydrodynamical instability grows only momentarily at each pericenter passage, an eventual runaway in the mode amplitude requires multiple orbits to occur. On the other hand, this accumulation allows runaway to occur at pericenter separations much greater than the prediction of Fig. 2. For the non-rotating case, the left panel of Fig. 4 shows that the runaway can kick in at a separation as large as $D_{p}/R_{t}\simeq 3.20$, whereas without the dynamical tide, the separation needs to be $D_{p}\lesssim 2.93R_{t}$ based on Fig. 2. Rotation suppresses the dynamical tide slightly (D. Lai, 1997), requiring a smaller $D_{p}/R_{t}\simeq 3.05$ for the runaway to occur within 20 orbits. We also note a larger variation in $|q_{2}|$ away from the pericenter. This is due to the beat of the $c_{2,+}$ (prograde) and $c_{-2,+}$ (retrograde) modes entering $q_{2}$. Note from Eq. (16) that the two modes have different natural frequencies when the background is rotating.
@@
Note the mode evolution largely resembles the chaotic tide studied in, e.g., R. A. Mardling (1995); P. B. Ivanov & J. C. B. Papaloizou (2004, 2007); Y. Wu (2018); M. Vick & D. Lai (2018); H. Yu et al. (2021, 2022). We will elaborate on this point further in §3.1.
@@
Figure 5: Numerical evolution of the $q_{2}$ mode for a non-rotating donor with $(n,\Gamma)=(1,2)$ in an eccentric orbit with $e=0.9$. Both the exact (left) and three-wave (right) nonlinear equations are considered. Runaways in mode amplitudes are observed when $D_{p}/R_{t}\lesssim 2.7$, matching nicely the predictions of J. Guillochon et al. (2011). The threshold $D_{p}/R_{t}$ can be even greater if the system is integrated for more orbits.
@@
As our theoretical study is largely motivated by the numerical simulations of J. Guillochon et al. (2011), we consider in Fig. 5 the evolution of a donor with $n=1$ and $\Gamma=2$ (a decent approximation for gaseous giant planets) moving in an eccentric orbit with $e=0.9$. In the affine model, the linear tidal overlap ($\propto I_{m}$) of an $n>0$ model is smaller than that of the $n=0$ model by a factor of $(5-n)\mathcal{M}_{\ast}$ with $\mathcal{M}_{\ast}=(\int x^{2}dM)/3$; see Appendix A. For $n=1$, $(5-n)\mathcal{M}_{\ast}\simeq 0.523$. As a result, the $n=1$ model requires a $D_{p}/R_{t}$ smaller than the $n=0$ model by a factor of $\sim 1/[(5-n)\mathcal{M}_{\ast}]^{1/3}\simeq 1.24$ for the runaway to occur. Other than this overall scaling difference, the evolution of the $n=1$ model largely resembles the $n=0$ model, and the predictions of the exact nonlinear equations (left) can be largely captured by the three-wave equations (right). Remarkably, our results show nice consistency with the simulations of J. Guillochon et al. (2011). Although the donor is not significantly disrupted in a single passage, the perturbation accumulates chaotically, eventually leading to a runaway in the mode amplitude in $\mathcal{O}(10)$ orbits when $D_{p}/R_{t}\lesssim 2.7R_{t}$.
@@
In Fig. 6, we show a survey of the maximum change in the mode energy $\Delta E_{\rm mode}$ reached in 22 orbits versus the orbital eccentricity $e$ and pericenter separation $D_{p}/R_{t}$. The mode energy is defined as $\Delta E_{\rm mode}=\Delta T+\Delta U+\Delta\Omega$, corresponding to the sum of the kinetic energy $\Delta T$ (Eq. A7), internal energy $\Delta U$ (Eq. A8), and self-gravity $\Delta\Omega$ (Eq. A18). At the leading order and for $n=0$, $\Delta E_{\rm mode}\simeq(1/2)\sum_{m}(\dot{q}_{m}^{2}/\omega_{f}^{2}+|q_{m}|^{2})$ in our normalization. The exact nonlinear equations from the affine model are used in the hydrodynamical evolution and the energy calculation (see Appendix A; B. Carter & J. P. Luminet 1985; P. Diener et al. 1995). The donor is assumed to be non-rotating. In the left and right panels, we use respectively $(n,\Gamma)=(0,20)$ and $(n,\Gamma)=(1,2)$. In the color scale, we saturate at values where runaways in the mode amplitudes are observed. We empirically determine these saturation values by plotting $\Delta E_{\rm mode}$ as a function of $|q_{2}|$, which leads to two clearly separate sectors. For smaller values $|q_{2}|$, we observe a nearly quadratic relation $\Delta E\propto|q_{2}|^{2}$, corresponding to bounded oscillations. The other sector has $|q_{2}|$ piled at unity (a criterion we use to terminate the integration), and at this fixed $|q_{2}|$, $\Delta E_{\rm mode}$ is multi-valued. The maximum $\Delta E_{\rm mode}$ from the first branch is taken to set the maximum in the color bar.
We note an approximately linear increase in the threshold $D_{p}/R_{t}$ with increasing orbital eccentricity $e$. This is the most critical modification of the earlier prediction of J. F. Sepinsky et al. (2007), which mainly considered the equipotential and found an expanded Roche lobe in eccentric orbits (harder for mass loss to occur; opposite of our results). Our study, therefore, highlights the significance of accounting for the hydrodynamical response of the fluid. Compared to the prediction assuming instantaneous equilibrium tide in Fig. 2, the accumulation of the dynamical tide over multiple orbits makes the runaways occur at even greater values of $D_{p}/R_{t}$. We also note that at fixed $e$, the energy $\Delta E_{\rm mode}$ does not vary monotonically from red to blue as $D_{p}$ increases. Instead, a random alternation of red and blue is observed near the threshold, indicating the chaotic nature of the evolution (see §3.1 for more discussions). We further emphasize that Fig. 6 serves only as a conservative estimation of the Roche limit, as the integration lasts only 22 orbits. Figs. 4 and 5 suggest that more systems may run away given more orbits, as the evolution is chaotic. However, integrating the system for longer would require a more careful treatment of the numerical scheme beyond brute-force integration of the differential equations to improve both efficiency and accuracy. This is left for future investigations.
@@
So far, the discussions have ignored damping. If there exists some efficient damping mechanism that could remove all the energy in the dynamical tide within one orbit, we would need to consider the hydrodynamical stability in a single passage. Fig. 7 shows the energy change acquired after a single passage (still assuming non-rotating $M$). When plotting $\Delta E_{\rm mode}$ versus $|q_{2}|$, we do not see a clear split into two sectors, and $|q_{2}|$ does not grow exponentially. Instead, at locations where bounded oscillations terminate in the multi-orbit case, we note $\Delta E_{\rm mode}$ becomes multi-valued at a given $|q_{2}|$. Such branching suggests instability to additional perturbations (and runaways occur in the next orbit), so we still saturate the color bar as before. We’ve also checked with three-wave equations of motion (which predict runaways in a single passage) and found consistent thresholds in $D_{p}/R_{t}$ as in Fig. 7 (which also approximately match P. Diener et al. 1995).
@@
This time, we note a decrease in $\Delta E_{\rm mode}$ with eccentricity. This is because the tidal interaction is significant only for a fraction of $2\pi/(\dot{\Phi}_{p}P_{\rm orb})\propto\sqrt{(1-e)^{3}/(1+e)}$ of the orbit. Although the mode becomes instantaneously unstable near the pericenter as long as $D_{p}/R_{t}$ is below the prediction of Fig. 2, the time for the instability to grow can be too short compared to the instability growth timescale, thereby preventing a runaway that could lead to mass loss. The threshold $D_{p}/R_{t}$ can differ by order unity when comparing Fig. 6 and 7 due to the dynamical tide. Using the linear damping calculation of P. Goldreich & P. D. Nicholson (1977) with a quality factor of $\mathcal{O}(10^{13})$, it requires $1-e\lesssim 10^{-8}$ to suppress the dynamical tide within a single orbit. Even if the quality factor is reduced to $10^{6}$ (Y. Wu, 2005), damping is significant only when $1-e<10^{-3}$. For systems less eccentric, Fig. 6 should be a more accurate representation of the Roche limit.
3 Applications
@@
After describing the general theoretical results in §2, we now describe the applications of the model in various astrophysical settings.
3.1 Relation to the chaotic tide
@@
It is interesting to note that the numerical mode evolution studied in §2.3 largely resembles the chaotic tide studied in, e.g., R. A. Mardling (1995); P. B. Ivanov & J. C. B. Papaloizou (2004, 2007); Y. Wu (2018); M. Vick & D. Lai (2018); H. Yu et al. (2021, 2022); S. Y. Lau & H. Yu (2025). In fact, our work presents a new mechanism to trigger the chaotic tide. In all previous analyses, the chaotic behavior ultimately originates from a random phase accumulation away from the pericenter. Such a random phase could be produced by fluctuations either in the orbital periods due to tidal back-reactions (R. A. Mardling, 1995; P. B. Ivanov & J. C. B. Papaloizou, 2004; Y. Wu, 2018; M. Vick & D. Lai, 2018) or gravitational-wave radiation (P. B. Ivanov & J. C. B. Papaloizou, 2007; S. Y. Lau & H. Yu, 2025), or in the mode frequencies due to nonlinear anharmonicities (H. Yu et al., 2021, 2022). A limitation of this mechanism is that the chaotic evolution can hardly be maintained once the orbital eccentricity drops below $e\simeq 0.9$ and the orbital period is not long enough for the random phase to accumulate. For example, for $M^{\prime}=10^{3}$, $e=0.7$, and $D_{p}/R_{t}=2.645$, we have $P_{\rm orb}=164.5$ (deliberately chosen to be far away from linear resonance) and $E_{\rm orb}=-MM^{\prime}(1-e)/(2D_{p})=-5.7$ (all in natural units). After a single passage, the mode acquires an energy $\Delta E_{\rm mode}\simeq 1.7\times 10^{-3}$ for the $(n,\Gamma)=(1,2)$ model. The one-pass perturbation of the mode phase from the tidal back-reaction is $2\Delta\phi_{q}=\omega_{f}\Delta P_{\rm orb}\simeq(-3/2)\omega_{f}P_{\rm orb}(\Delta E_{\rm mode}/E_{\rm orb})\simeq 0.091\,{\rm rad}$, much smaller than the threshold value of $2\Delta\phi_{q}\gtrsim 1\,{\rm rad}$ derived in M. Vick & D. Lai (2018). Moreover, back-reaction is turned off in our simulations. The anharmonicity-induced phase perturbation differs from the back-reaction phase by a factor of $\mathcal{O}(|E_{\rm orb}|)$ and typically has a greater numerical value (H. Yu et al., 2021), yet $2\Delta\phi_{q}=P_{\rm orb}\Delta\omega_{f}\sim\omega_{f}P_{\rm orb}\Delta E_{\rm mode}\simeq 0.34\,{\rm rad}$ is still smaller than unity to trigger chaos. Nonetheless, our Fig. 6 shows that this system indeed grows chaotically and eventually runs away.
@@
Crucially, all studies referenced above used the linear theory of W. H. Press & S. A. Teukolsky (1977); D. Lai (1997) to compute the tidal kick each mode receives at the pericenter. In this work, we incorporate nonlinear hydrodynamics at each pericenter passage, which may allow the mode to grow momentarily due to instability. If the mode’s phase at pericenter is periodic, then the growth will repeat after a few orbits and accumulate. This allows the mode to reach an amplitude large enough to trigger chaos. The growth then continues chaotically until the mode eventually runs away.

### v2

@@ II.3 Dynamical tide induced instability and diffusive tide
For binaries with significant eccentricities, the truly dynamical component of a mode can be excited at a pericenter passage. Such a component oscillates at a mode’s natural frequency (hence phase incoherent with the orbit) and does not decrease in amplitude as the companion moves to the apocenter (assuming dissipation is small). In the next pericenter passage, the existing component of the dynamical tide can interfere with the new tidal kick it receives at the pericenter, hence making this component’s evolution history dependent (as opposed to the equilibrium tide that depends only on the instantaneous orbital configuration). This dynamical component formally modifies the stability threshold according to the toy model in Eq. (7), though its complicated evolution prohibits a simple analytical study as we have done in §II.1 and II.2. Nonetheless, the hydrodynamical equations are sufficiently generic for arbitrary rotations and orbital configurations. We will use a numerical approach to investigate the associated hydrodynamical stability.
@@
In this Subsection, we will solve directly the phase-space expanded equations of $c_{m,\pm}$, Eqs. (16), so that the Coriolis effect can be accounted for naturally. For the driving terms ($K_{m}$), we will consider both the version expanded to the three-wave order, Eqs. (21) and (22), and the exact nonlinear version derived in Appendix A following the affine model of Diener et al. (1995) (see also Carter and Luminet 1983, 1985). The monopolar mode $q_{\rm tr}$ is treated fully dynamically, where a differential equation of $\ddot{q}_{\rm tr}$ is solved instead of reducing $q_{\rm tr}$ under the static limit with Eqs. (A31). Note that the affine model in Diener et al. (1995) allows for generic values of polytropic index $n$ and adiabatic index $\Gamma$. Thus, our numerical experiments will consider both a homogeneous ellipsoid with $(n,\ \Gamma)=(0,\ 20)$ as before [^3] Numerically, $\Gamma\to\infty$ is hard to implement. We therefore use a finite $\Gamma=20\gg 1$ as an approximation. , and a polytrope with $(n,\ \Gamma)=(1,\ 2)$ that serves as a good approximation of a buoyancy-neutral Jovian planet. However, a caveat for the $n=1$ model is that the starting point of the affine model (Eq. 13) is not self-consistently satisfied as the Lagrangian displacement $\boldsymbol{\xi}$ is no longer linear in position. We fix the companion’s mass to be $M^{\prime}=10^{3}$ and treat it as a point particle. When computing the orbital motion, we ignore tidal back-reaction and simply model it as Keplerian, with
@@
The initial conditions are set at the apocenter, with the modes computed with their linear equilibrium tide solution. The set of differential equations is evolved using the solve_ivp function in SciPy (Virtanen et al., 2020) with an explicit Runge-Kutta method of order 8 (Hairer et al., 1993).
@@
In the first experiment shown in Fig. 3, we validate the results of §II.2 by considering the numerical evolution of $|q_{2}|=|c_{2,+}+c_{-2,+}^{\ast}|$ for a non-rotating ($p=0$), homogeneous ($n=0$) star in a moderately eccentric orbit with $e=0.3$. We evolve the system using both the exact nonlinear equations (left panel) and the three-wave-expanded ones (right panel). Qualitatively, the two sets of equations yield similar results, with the exact version allowing the runaway to occur at a slightly further separation $D_{p}/R_{t}\simeq 2.72$. This is consistent with Fig. 1, which shows that the three-wave equations slightly underestimates $D_{p}^{(\rm th)}\propto(1/\epsilon_{\rm max})^{1/3}$.
@@
The accumulation of the dynamical tide becomes more significant in systems with higher eccentricities, with one such example shown in Fig. 4 for a donor moving in a highly eccentric ($e=0.9$) orbit. We still consider a homogeneous donor but vary its rotation rates (left for the non-rotating case and right for the pseudo-synchronized case). Both cases are evolved using the three-wave equations of motion. As the hydrodynamical instability grows only momentarily at each pericenter passage, an eventual runaway in the mode amplitude requires multiple orbits to occur. On the other hand, this accumulation allows runaway to occur at pericenter separations much greater than the prediction of Fig. 2. For the non-rotating case, the left panel of Fig. 4 shows that the runaway can kick in at a separation as large as $D_{p}/R_{t}\simeq 3.20$, whereas without the dynamical tide, the separation needs to be $D_{p}\lesssim 2.93R_{t}$ based on Fig. 2. Rotation suppresses the dynamical tide slightly (Lai, 1997), requiring a smaller $D_{p}/R_{t}\simeq 3.05$ for the runaway to occur within 20 orbits. We also note a larger variation in $|q_{2}|$ away from the pericenter. This is due to the beat of the $c_{2,+}$ (prograde) and $c_{-2,+}$ (retrograde) modes entering $q_{2}$. Note from Eq. (17) that the two modes have different natural frequencies when the background is rotating.
@@
Note the mode evolution largely resembles the diffusive tide (also known as the chaotic tide) studied in, e.g., Mardling (1995); Ivanov and Papaloizou (2004, 2007); Wu (2018); Vick and Lai (2018); Yu et al. (2021, 2022). We will elaborate on this point further in §III.1.
@@
Figure 5: Numerical evolution of the $q_{2}$ mode for a non-rotating donor with $(n,\Gamma)=(1,2)$ in an eccentric orbit with $e=0.9$. Both the exact (left) and three-wave (right) nonlinear equations are considered. Runaways in mode amplitudes are observed when $D_{p}/R_{t}\lesssim 2.7$, matching nicely the predictions of Guillochon et al. (2011). The threshold $D_{p}/R_{t}$ can be even greater if the system is integrated for more orbits.
@@
As our theoretical study is largely motivated by the numerical simulations of Guillochon et al. (2011), we consider in Fig. 5 the evolution of a donor with $n=1$ and $\Gamma=2$ (a decent approximation for gaseous giant planets) moving in an eccentric orbit with $e=0.9$. In the affine model, the linear tidal overlap ($\propto I_{m}$) of an $n>0$ model is smaller than that of the $n=0$ model by a factor of $(5-n)\mathcal{M}_{\ast}$ with $\mathcal{M}_{\ast}=(\int x^{2}dM)/3$; see Appendix A. For $n=1$, $(5-n)\mathcal{M}_{\ast}\simeq 0.523$. As a result, the $n=1$ model requires a $D_{p}/R_{t}$ smaller than the $n=0$ model by a factor of $\sim 1/[(5-n)\mathcal{M}_{\ast}]^{1/3}\simeq 1.24$ for the runaway to occur. Other than this overall scaling difference, the evolution of the $n=1$ model largely resembles the $n=0$ model, and the predictions of the exact nonlinear equations (left) can be largely captured by the three-wave equations (right). Remarkably, our results show nice consistency with the simulations of Guillochon et al. (2011). Although the donor is not significantly disrupted in a single passage, the perturbation accumulates diffusively, eventually leading to a runaway in the mode amplitude in $\mathcal{O}(10)$ orbits when $D_{p}/R_{t}\lesssim 2.7R_{t}$.
@@
In Fig. 6, we show a survey of the maximum change in the mode energy $\Delta E_{\rm mode}$ reached in 22 orbits versus the orbital eccentricity $e$ and pericenter separation $D_{p}/R_{t}$. The mode energy is defined as $\Delta E_{\rm mode}=\Delta T+\Delta U+\Delta\Omega$, corresponding to the sum of the kinetic energy $\Delta T$ (Eq. A7), internal energy $\Delta U$ (Eq. A9), and self-gravity $\Delta\Omega$ (Eq. A18). At the leading order and for $n=0$, $\Delta E_{\rm mode}\simeq(1/2)\sum_{m}(\dot{q}_{m}^{2}/\omega_{f}^{2}+|q_{m}|^{2})$ in our normalization. The exact nonlinear equations from the affine model are used in the hydrodynamical evolution and the energy calculation (see Appendix A; Carter and Luminet 1985; Diener et al. 1995). The donor is assumed to be non-rotating. In the left and right panels, we use respectively $(n,\Gamma)=(0,20)$ and $(n,\Gamma)=(1,2)$. In the color scale, we saturate at values where runaways in the mode amplitudes are observed. We empirically determine these saturation values by plotting $\Delta E_{\rm mode}$ as a function of $|q_{2}|$, which leads to two clearly separate sectors. For smaller values $|q_{2}|$, we observe a nearly quadratic relation $\Delta E\propto|q_{2}|^{2}$, corresponding to bounded oscillations. The other sector has $|q_{2}|$ piled at unity (a criterion we use to terminate the integration), and at this fixed $|q_{2}|$, $\Delta E_{\rm mode}$ is multi-valued. The maximum $\Delta E_{\rm mode}$ from the first branch is taken to set the maximum in the color bar.
We note an approximately linear increase in the threshold $D_{p}/R_{t}$ with increasing orbital eccentricity $e$. This is the most critical modification of the earlier prediction of Sepinsky et al. (2007), which mainly considered the equipotential and found an expanded Roche lobe in eccentric orbits (harder for mass loss to occur; opposite of our results). Our study, therefore, highlights the significance of accounting for the hydrodynamical response of the fluid. Compared to the prediction assuming instantaneous equilibrium tide in Fig. 2, the accumulation of the dynamical tide over multiple orbits makes the runaways occur at even greater values of $D_{p}/R_{t}$. We also note that at fixed $e$, the energy $\Delta E_{\rm mode}$ does not vary monotonically from red to blue as $D_{p}$ increases. Instead, a random alternation of red and blue is observed near the threshold, indicating the stochastic nature of the evolution (see §III.1 for more discussions). We further emphasize that Fig. 6 serves only as a conservative estimation of the Roche limit, as the integration lasts only 22 orbits. Figs. 4 and 5 suggest that more systems may run away given more orbits, as the evolution is diffusive. However, integrating the system for longer would require a more careful treatment of the numerical scheme beyond brute-force integration of the differential equations to improve both efficiency and accuracy. This is left for future investigations.
@@
So far, the discussions have ignored damping. If there exists some efficient damping mechanism that could remove all the energy in the dynamical tide within one orbit, we would need to consider the hydrodynamical stability in a single passage. Fig. 7 shows the energy change acquired after a single passage (still assuming non-rotating $M$). When plotting $\Delta E_{\rm mode}$ versus $|q_{2}|$, we do not see a clear split into two sectors, and $|q_{2}|$ does not grow exponentially. Instead, at locations where bounded oscillations terminate in the multi-orbit case, we note $\Delta E_{\rm mode}$ becomes multi-valued at a given $|q_{2}|$. Such branching suggests instability to additional perturbations (and runaways occur in the next orbit), so we still saturate the color bar as before. We’ve also checked with three-wave equations of motion (which predict runaways in a single passage) and found consistent thresholds in $D_{p}/R_{t}$ as in Fig. 7 (which also approximately match Diener et al. 1995).
@@
This time, we note a decrease in $\Delta E_{\rm mode}$ with eccentricity. This is because the tidal interaction is significant only for a fraction of $2\pi/(\dot{\Phi}_{p}P_{\rm orb})\propto\sqrt{(1-e)^{3}/(1+e)}$ of the orbit. Although the mode becomes instantaneously unstable near the pericenter as long as $D_{p}/R_{t}$ is below the prediction of Fig. 2, the time for the instability to grow can be too short compared to the instability growth timescale, thereby preventing a runaway that could lead to mass loss. The threshold $D_{p}/R_{t}$ can differ by order unity when comparing Fig. 6 and 7 due to the dynamical tide. Using the linear damping calculation of Goldreich and Nicholson (1977) with a quality factor of $\mathcal{O}(10^{13})$, it requires $1-e\lesssim 10^{-8}$ to suppress the dynamical tide within a single orbit. Even if the quality factor is reduced to $10^{6}$ (Wu, 2005), damping is significant only when $1-e<10^{-3}$. For systems less eccentric, Fig. 6 should be a more accurate representation of the Roche limit. As a caveat, the discussion considers only linear damping. Kumar and Goodman (1996) suggested that nonlinear damping can be highly efficient with a quality factor of $\mathcal{O}(100)$. However, their calculation focused on main sequence stars with masses $0.5-0.75\,M_{\odot}$ and was based on order of magnitude estimations. Whether nonlinear damping can be as strong as claimed in Kumar and Goodman (1996) for donors of different types remain to be examined in details by future studies.
III Applications
@@
After describing the general theoretical results in §II, we now describe the applications of the model in various astrophysical settings.
III.1 Relation to the diffusive tide
@@
It is interesting to note that the numerical mode evolution studied in §II.3 largely resembles the diffusive (or chaotic) tide studied in, e.g., Mardling (1995); Ivanov and Papaloizou (2004, 2007); Wu (2018); Vick and Lai (2018); Yu et al. (2021, 2022); Lau and Yu (2025). In fact, our work presents a new mechanism to trigger the diffusive tide. In all previous analyses, the diffusive behavior ultimately originates from a random phase accumulation away from the pericenter. Such a random phase could be produced by fluctuations either in the orbital periods due to tidal back-reactions (Mardling, 1995; Ivanov and Papaloizou, 2004; Wu, 2018; Vick and Lai, 2018) or gravitational-wave radiation (Ivanov and Papaloizou, 2007; Lau and Yu, 2025), or in the mode frequencies due to nonlinear anharmonicities (Yu et al., 2021, 2022). A limitation of this mechanism is that the diffusive evolution can hardly be maintained once the orbital eccentricity drops below $e\simeq 0.9$ and the orbital period is not long enough for the random phase to accumulate. For example, for $M^{\prime}=10^{3}$, $e=0.7$, and $D_{p}/R_{t}=2.645$, we have $P_{\rm orb}=164.5$ (deliberately chosen to be far away from linear resonance) and $E_{\rm orb}=-MM^{\prime}(1-e)/(2D_{p})=-5.7$ (all in natural units). After a single passage, the mode acquires an energy $\Delta E_{\rm mode}\simeq 1.7\times 10^{-3}$ for the $(n,\Gamma)=(1,2)$ model. The one-pass perturbation of the mode phase from the tidal back-reaction is $2\Delta\phi_{q}=\omega_{f}\Delta P_{\rm orb}\simeq(-3/2)\omega_{f}P_{\rm orb}(\Delta E_{\rm mode}/E_{\rm orb})\simeq 0.091\,{\rm rad}$, much smaller than the threshold value of $2\Delta\phi_{q}\gtrsim 1\,{\rm rad}$ derived in Vick and Lai (2018). Moreover, back-reaction is turned off in our simulations. The anharmonicity-induced phase perturbation differs from the back-reaction phase by a factor of $\mathcal{O}(|E_{\rm orb}|)$ and typically has a greater numerical value (Yu et al., 2021), yet $2\Delta\phi_{q}=P_{\rm orb}\Delta\omega_{f}\sim\omega_{f}P_{\rm orb}\Delta E_{\rm mode}\simeq 0.34\,{\rm rad}$ is still smaller than unity to trigger the diffusive growth. Nonetheless, our Fig. 6 shows that this system indeed grows diffusively and eventually runs away.
@@
Crucially, all studies referenced above used the linear theory of Press and Teukolsky (1977); Lai (1997) to compute the tidal kick each mode receives at the pericenter. In this work, we incorporate nonlinear hydrodynamics at each pericenter passage, which may allow the mode to grow momentarily due to instability. If the mode’s phase at pericenter is periodic, then the growth will repeat after a few orbits and accumulate. This allows the mode to reach an amplitude large enough to trigger the diffusive tide. The growth then continues stochastically until the mode eventually runs away.